# Quantum conditional mutual information as a probe of measurement-induced entanglement phase transitions

### v1

@@ III.3 Data-collapse
Here, we adopt an empirical approach to estimate $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit based on a data-collapse analysis. We employ the Bayesian scaling analysis method [35] to collapse the data of $I_{\gamma}(p,L)$, assuming the following scaling ansatz:
@@
where $\Delta p=p-p_{\mathrm{c}}$ is the deviation from the critical point, and $f_{\gamma}$ denotes the scaling function. Representative examples of successful data-collapse fits are shown in Fig. 5. We do not explicitly include a correction term of the form $\lambda L^{-\omega}$ in the scaling function, where $\lambda$ is an irrelevant scaling field and $\omega$ is the associated correction exponent. Instead, we perform data-collapse fits over the system size range $L\in[L_{\mathrm{min}},L_{\mathrm{max}}]$ with $L_{\mathrm{max}}=2048$ fixed and $L_{\mathrm{min}}$ systematically varied. We then examine the asymptotic behavior of the estimated values of $p_{\mathrm{c}}$ and $\nu$ as $L_{\mathrm{min}}$ increases [36, 37]. An important aspect of this analysis is the estimation of uncertainties in the fitted parameters. Because data collapsing is inherently a nonlinear procedure, the results are sensitive to the choice of initial parameters in the Bayesian analysis as well as the selected data set. To mitigate these sensitivities and obtain reliable error bars, we apply the bootstrap resampling method. The resulting uncertainties of the estimated $p_{\mathrm{c}}$ and $\nu$ are shown in Figs. 6 and 7. We extrapolate the results to the thermodynamic limit, i.e., $01/L_{\mathrm{min}}=$ 0, using a linear fit based on the four largest values of $L_{\mathrm{min}}$. Although the use of a linear fit and the specific number of data points are empirical choices, we find that the estimated values of $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit are consistent between $\gamma=a$ and $\gamma=b$ within two standard deviations. As summarized in Table 1, the results obtained for other values of $\alpha$ are overall consistent between the two different partitioning schemes, agreeing to at least four (three) significant digits for $p_{\mathrm{c}}$ ($\nu$) in most cases of $\alpha$. Additional analyses are provided in Supplemental Material [23].
@@
Figure 6: Extrapolation of (a) $p_{\mathrm{c}}$ and (b) $\nu$ obtained from data-collapse analysis using the QCMI $I_{\gamma}(p,L)$ for $L\geqslant L_{\mathrm{min}}$ at $\alpha=3.5$. Solid and dashed lines represent linear fits to the data for $\gamma=a$ and $\gamma=b$, respectively. The extrapolated values in the thermodynamic limit ($1/L_{\mathrm{min}}\to 0$) are indicated by solid symbols at $01/L_{\mathrm{min}}=$ 0.
@@ III.4 Crossing-point analysis
With the accurate estimates of $p_{\mathrm{c}}$ and $\nu$ obtained from the data-collapse analysis, we now return to the crossing-point analysis. Crossing points are defined as the values of $p$ and $I_{\gamma}$ at which the curves of $I_{\gamma}(p,L)$ and $I_{\gamma}(p,2L)$ intersect. These points are denoted as $p_{\gamma}^{\times}(L,2L)$ and $I_{\gamma}^{\times}(L,2L)$, respectively. A key difficult in this analysis is the nonmonotonic finite-size behavior of the crossing points, as clearly illustrated in Figs. 8(a) and 9(a). To capture this behavior, it is necessary to consider at least two correction terms in the scaling forms:
@@
where $a_{i}$ and $b_{i}$ ($i$=1,2) are constants, and the correction exponents are ordered such that $\omega_{2}>\omega_{1}>0$. Note that $I_{b}^{\times}(L,2L)$ is normalized by a constant factor ($\simeq 1.83$) so that its thermodynamic limit matches that of $I_{a}^{\times}(L,2L)$, i.e., $c/3$, [see Eqs. (7) and (8)]. There are two possible scenarios for the origin of the second correction term [34]. One possibility is that it arises from a second irrelevant scaling field, independent of the first. The other is that it corresponds to the second higher-order term of the same irrelevant filed as the first one, i.e, $\omega_{2}=2\omega_{1}$. Although the scaling forms in Eqs. (10) and (11) appear relatively simple, we find that fitting both correction terms with sufficient precision is quite challenging [34, 33]. This difficulty persists even when we fix $p_{\mathrm{c}}$ and $\nu$ in Eq. (10) to the values obtained from the data-collapse analysis in Table 1.
@@
Figure 8: System-size dependence of the crossing points $p_{\gamma}^{\times}(L,2L)$ for (a) $\alpha=3.5$ and (b) $\alpha=2.5$. Open circles (triangles) represent the results for $\gamma=a$ ($b$), while solid ones at $01/L=$ 0 indicate the values of $p_{\mathrm{c}}$ obtained by the data-collapse analysis shown in Figs. 6 and 7, as well as in Table 1. Solid and dashed lines are independent fits to the data for the two partitioning schemes, $\gamma=a$ and $\gamma=b$, respectively, using the functional from given in Eq. (10), where $p_{\mathrm{c}}$ and $\nu$ are fixed at the extrapolated values in the thermodynamic limit obtained by the data-collapse analysis, and $\omega_{2}=2\omega_{1}$ is assumed. The fitting values of $\omega_{1}$ are indicated in each panel.
@@ IV Discussion
In Fig. 10, we plot the correlation-length exponent $\nu$ estimated from the data-collapse analysis (see Table 1) as a function of $\alpha$. We confirm that for $\alpha\gtrsim 3$, our model belongs to the same universality class as that of the short-range interacting models [13, 15]. In particular, our numerical estimates are in full agreement with recent results for the model with nearest-neighbor Clifford gates [19], and differ significantly from the value predicted by classical two-dimensional percolation theory, $\nu=4/3$.
The existence of a universality class behavior for $\alpha\gtrsim 3$ is further supported by the coefficient $c$ of the logarithmic term in the entanglement entropy, as shown in Fig. 11. The close agreement of $c$ values estimated from the two different partitioning schemes, $\gamma=a$ and $b$, strongly supports the presence of emergent conformal criticality as described by Eqs. (5)–(8) for $\alpha\gtrsim 3$, in line with the behavior observed in the short-range interacting models. In contrast, the noticeable discrepancy in the estimated values of $c$ between the two partitioning schemes for $\alpha\lesssim 3$ indicates a breakdown of the scaling relations in Eqs. (5)–(8). This deviation from emergent conformal criticality corroborates previous findings of non-conformal-field-theory criticality, as evidenced by the deviation of the dynamical critical exponent $z$ from unity in Ref. [13]. It is worth noting, however, that the critical point $p_{c}$ and the exponent $\nu$ remain consistent between the two partitioning schemes for all values of $\alpha$ (see Table 1). This consistency highlights that the QCMI remains a valid and robust probe of the measurement-induced entanglement phase transition, even in the absence of emergent conformal criticality.
Figure 10: $\alpha$-dependence of the correlation-length critical exponent $\nu$. The dotted line with the gray shaded area represents the result of the nearest-neighbor model, $\nu=1.265(15)$ [19]. The inset shows a magnified view near $\alpha=3.0$. For comparison, the prediction of classical two-dimensional percolation theory, $\nu=4/3$, is also indicated by the dashed line.
Now we compare the coefficient $c/3$ of the logarithmic term in the entanglement entropy with those obtained for monitored nearest-neighbor random Clifford circuits in Refs. [2, 11, 19]. While our analysis uses the base-2 logarithm, previous studies employed the natural logarithm in defining this coefficient. To enable direct comparison, we convert our result as
@@
so that $\tilde{c}$ is consistent with the convention used in the earlier works. From our estimate of $c/3=1.053(2)$ for the shortest-range case $\alpha=4$ using the partitioning scheme $\gamma=a$ (see Table 2), we obtain $\tilde{c}=1.519(3)$. This value is noticeably smaller than those in the previous studies: $\tilde{c}=1.6$ with up to 512 qubits [2], $\tilde{c}=1.61(3)$ with up to 512 qubits [11], and $\tilde{c}=1.57(1)$ with up to 10240 qubits [19] for monitored nearest-neighbor random Clifford circuits. Furthermore, we find that $c/3$ exhibits only minor variation across different partitioning schemes ($\gamma=a,b$) and for a range of $\alpha\gtrsim 3$ [see the inset of Fig. 11 and Table 2]. This supports the conclusion that our estimate of $\tilde{c}$ remains consistently smaller than those in the previous works, even when accounting for uncertainties.
@@
To clarify the origin of this discrepancy, we also estimate $\tilde{c}$ from the entanglement entropy at criticality, as shown in Fig. 12. Using the critical point $p_{\mathrm{c}}$ obtained from the data-collapse analysis, we fit the entanglement entropy of the form given in Eq. (5), treating $c/3$ and $c^{\prime}$ as fitting parameters. We obtain $\tilde{c}=1.6126(7)$, $1.5484(9)$, and $1.5263(7)$ [$c/3=1.1175(5)$, $1.0730(6)$, and $1.0577(5)$] for $\alpha=3.0$, $3.5$, and $4.0$, respectively. These results indicate that $c/3$ obtained from the entanglement entropy approaches the value estimated from the QCMI from above as $\alpha$ increases (see Supplemental Material [23] for results at other values of $\alpha$). Since finite-size effects are expected to be smaller at larger $\alpha$, the estimates of $\tilde{c}$ from the entanglement entropy at criticality are more reliable for larger $\alpha$ than for the smaller ones. Nevertheless, we argue that the estimates of $\tilde{c}$ based on the QCMI are generally more accurate, as the QCMI enables crossing-point analyses by eliminating the logarithmic dependence at criticality, as shown in Eqs. (7) and (8). Indeed, the values of $\tilde{c}$ estimated via the QCMI exhibit much weaker dependence on $\alpha$ [see the inset of Fig. 11 and Table 2] compared to those obtained from the entanglement entropy [23]. We therefore adopt $\tilde{c}=1.519(3)$, obtained from the QCMI for $\alpha=4.0$, as our final estimate for the shortest-range interacting case.

### v2

@@ III.3 Data-collapse
Here, we adopt an empirical approach to estimate $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit based on a data-collapse analysis. We employ the Bayesian scaling analysis method [Harada-PRE2011] to collapse the data of $I_{\gamma}(p,L)$, assuming the following scaling ansatz:
@@
where $\Delta p=p-p_{\mathrm{c}}$ is the deviation from the critical point, and $f_{\gamma}$ denotes the scaling function. Representative examples of successful data-collapse fits are shown in Fig. 5. We do not explicitly include a correction term of the form $\lambda L^{-\omega}$ in the scaling function, where $\lambda$ is an irrelevant scaling field and $\omega$ is the associated correction exponent. Instead, we perform data-collapse fits over the system size range $L\in[L_{\mathrm{min}},L_{\mathrm{max}}]$ with $L_{\mathrm{max}}=2048$ fixed and $L_{\mathrm{min}}$ systematically varied. We then examine the asymptotic behavior of the estimated values of $p_{\mathrm{c}}$ and $\nu$ as $L_{\mathrm{min}}$ increases [Otsuka-PRX2016, Otsuka-PRB2020]. An important aspect of this analysis is the estimation of uncertainties in the fitted parameters. Because data collapsing is inherently a nonlinear procedure, the results are sensitive to the choice of initial parameters in the Bayesian analysis as well as the selected data set. To mitigate these sensitivities and obtain reliable error bars, we apply the bootstrap resampling method. The resulting uncertainties of the estimated $p_{\mathrm{c}}$ and $\nu$ are shown in Figs. 6 and 7. We extrapolate the results to the thermodynamic limit, i.e., $1/L_{\mathrm{min}}=0$, using a linear fit based on the four largest values of $L_{\mathrm{min}}$. Although the use of a linear fit and the specific number of data points are empirical choices, we find that the estimated values of $p_{\mathrm{c}}$ and $\nu$ in the thermodynamic limit are consistent between $\gamma=a$ and $\gamma=b$ within two standard deviations. As summarized in Table 1, the results obtained for other values of $\alpha$ are overall consistent between the two different partitioning schemes, agreeing to at least four (three) significant digits for $p_{\mathrm{c}}$ ($\nu$) in most cases of $\alpha$. Additional analyses are provided in Supplemental Material [suppl].
@@
Figure 6: Extrapolation of (a) $p_{\mathrm{c}}$ and (b) $\nu$ obtained from data-collapse analysis using the QCMI $I_{\gamma}(p,L)$ for $L\geqslant L_{\mathrm{min}}$ at $\alpha=3.5$. Solid and dashed lines represent linear fits to the data for $\gamma=a$ and $\gamma=b$, respectively. The extrapolated values in the thermodynamic limit ($1/L_{\mathrm{min}}\to 0$) are indicated by solid symbols at $1/L_{\mathrm{min}}=0$.
@@ III.4 Crossing-point analysis
With the accurate estimates of $p_{\mathrm{c}}$ and $\nu$ obtained from the data-collapse analysis, we now return to the crossing-point analysis. Crossing points are defined as the values of $p$ and $I_{\gamma}$ at which the curves of $I_{\gamma}(p,L)$ and $I_{\gamma}(p,2L)$ intersect. These points are denoted as $p_{\gamma}^{\times}(L,2L)$ and $I_{\gamma}^{\times}(L,2L)$, respectively. A key difficulty in this analysis is the nonmonotonic finite-size behavior of the crossing points, as clearly illustrated in Figs. 8(a) and 9(a). To capture this behavior, it is necessary to consider at least two correction terms in the scaling forms:
@@
where $a_{i}$ and $b_{i}$ (ii=1,2) are constants, and the correction exponents are ordered such that $\omega_{2}>\omega_{1}>0$. Note that $I_{b}^{\times}(L,2L)$ is normalized by a constant factor ($\simeq 1.83$) so that its thermodynamic limit matches that of $I_{a}^{\times}(L,2L)$, i.e., $c/3$, [see Eqs. (7) and (8)]. There are two possible scenarios for the origin of the second correction term [Ma-PRL2018]. One possibility is that it arises from a second irrelevant scaling field, independent of the first. The other is that it corresponds to the second higher-order term of the same irrelevant filed as the first one, i.e, $\omega_{2}=2\omega_{1}$. Although the scaling forms in Eqs. (10) and (11) appear relatively simple, we find that fitting both correction terms with sufficient precision is quite challenging [Ma-PRL2018, Shao-S2016]. This difficulty persists even when we fix $p_{\mathrm{c}}$ and $\nu$ in Eq. (10) to the values obtained from the data-collapse analysis in Table 1.
@@
Figure 8: System-size dependence of the crossing points $p_{\gamma}^{\times}(L,2L)$ for (a) $\alpha=3.5$ and (b) $\alpha=2.5$. Open circles (triangles) represent the results for $\gamma=a$ ($b$), while solid ones at $1/L=0$ indicate the values of $p_{\mathrm{c}}$ obtained by the data-collapse analysis shown in Figs. 6 and 7, as well as in Table 1. Solid and dashed lines are independent fits to the data for the two partitioning schemes, $\gamma=a$ and $\gamma=b$, respectively, using the functional from given in Eq. (10), where $p_{\mathrm{c}}$ and $\nu$ are fixed at the extrapolated values in the thermodynamic limit obtained by the data-collapse analysis, and $\omega_{2}=2\omega_{1}$ is assumed. The fitting values of $\omega_{1}$ are indicated in each panel.
@@ IV Discussion
In Fig. 10, we plot the correlation-length exponent $\nu$ estimated from the data-collapse analysis (see Table 1) as a function of $\alpha$. We confirm that for $\alpha\gtrsim 3$, our model belongs to the same universality class as that of the short-range interacting models [Block-PRL2022, Sahu-PRB2022]. In particular, our numerical estimates are in full agreement with recent results for the model with nearest-neighbor Clifford gates [Sierant-PRB2022], and differ significantly from the value predicted by classical two-dimensional percolation theory, $\nu=4/3$.
The existence of a universality class behavior for $\alpha\gtrsim 3$ is further supported by the coefficient $c$ of the logarithmic term in the entanglement entropy, as shown in Fig. 11. The close agreement of $c$ values estimated from the two different partitioning schemes, $\gamma=a$ and $b$, strongly supports the presence of emergent conformal criticality as described by Eqs. (5)–(8) for $\alpha\gtrsim 3$, in line with the behavior observed in the short-range interacting models. In contrast, the noticeable discrepancy in the estimated values of $c$ between the two partitioning schemes for $\alpha\lesssim 3$ indicates a breakdown of the scaling relations in Eqs. (5)–(8). This deviation from emergent conformal criticality corroborates previous findings of non-conformal-field-theory criticality, as evidenced by the deviation of the dynamical critical exponent $z$ from unity in Ref. [Block-PRL2022]. It is worth noting, however, that the critical point $p_{c}$ and the exponent $\nu$ remain consistent between the two partitioning schemes for all values of $\alpha$ (see Table 1). This consistency highlights that the QCMI remains a valid and robust probe of the measurement-induced entanglement phase transition, even in the absence of emergent conformal criticality.
Figure 10: $\alpha$-dependence of the correlation-length critical exponent $\nu$. The dotted line with the gray shaded area represents the result of the nearest-neighbor model, $\nu=1.265(15)$ [Sierant-PRB2022]. The inset shows a magnified view near $\alpha=3.0$. For comparison, the prediction of classical two-dimensional percolation theory, $\nu=4/3$, is also indicated by the dashed line.
Now we compare the coefficient $c/3$ of the logarithmic term in the entanglement entropy with those obtained for monitored nearest-neighbor random Clifford circuits in Refs. [Li-PRB2019a, Zabalo-PRB2020, Sierant-PRB2022]. While our analysis uses the base-2 logarithm, previous studies employed the natural logarithm in defining this coefficient. To enable direct comparison, we convert our result as
@@
so that $\tilde{c}$ is consistent with the convention used in the earlier works. From our estimate of $c/3=1.053(2)$ for the shortest-range case $\alpha=4$ using the partitioning scheme $\gamma=a$ (see Table 2), we obtain $\tilde{c}=1.519(3)$. This value is noticeably smaller than those in the previous studies: $\tilde{c}=1.6$ with up to 512 qubits [Li-PRB2019a], $\tilde{c}=1.61(3)$ with up to 512 qubits [Zabalo-PRB2020], and $\tilde{c}=1.57(1)$ with up to 10240 qubits [Sierant-PRB2022] for monitored nearest-neighbor random Clifford circuits. Furthermore, we find that $c/3$ exhibits only minor variation across different partitioning schemes ($\gamma=a,b$) and for a range of $\alpha\gtrsim 3$ [see the inset of Fig. 11 and Table 2]. This supports the conclusion that our estimate of $\tilde{c}$ remains consistently smaller than those in the previous works, even when accounting for uncertainties.
@@
To clarify the origin of this discrepancy, we also estimate $\tilde{c}$ from the entanglement entropy at criticality, as shown in Fig. 12. Using the critical point $p_{\mathrm{c}}$ obtained from the data-collapse analysis, we fit the entanglement entropy of the form given in Eq. (5), treating $c/3$ and $c^{\prime}$ as fitting parameters. We obtain $\tilde{c}=1.6126(7)$, $1.5484(9)$, and $1.5263(7)$ [$c/3=1.1175(5)$, $1.0730(6)$, and $1.0577(5)$] for $\alpha=3.0$, $3.5$, and $4.0$, respectively. These results indicate that $c/3$ obtained from the entanglement entropy approaches the value estimated from the QCMI from above as $\alpha$ increases (see Supplemental Material [suppl] for results at other values of $\alpha$). Since finite-size effects are expected to be smaller at larger $\alpha$, the estimates of $\tilde{c}$ from the entanglement entropy at criticality are more reliable for larger $\alpha$ than for the smaller ones. Nevertheless, we argue that the estimates of $\tilde{c}$ based on the QCMI are generally more accurate, as the QCMI enables crossing-point analyses by eliminating the logarithmic dependence at criticality, as shown in Eqs. (7) and (8). Indeed, the values of $\tilde{c}$ estimated via the QCMI exhibit much weaker dependence on $\alpha$ [see the inset of Fig. 11 and Table 2] compared to those obtained from the entanglement entropy [suppl]. We therefore adopt $\tilde{c}=1.519(3)$, obtained from the QCMI for $\alpha=4.0$, as our final estimate for the shortest-range interacting case.